An economy's production function is given by: Y=2√K √AN , the labour force grows at a rate of 1% .The technology advances at rate of 6%. The depreciation and Saving rates are 0.10 and 0.20 respectively. What is the capital per effective workers in steady state?
1
Expert's answer
2021-09-27T15:44:40-0400

"Y = 2 \sqrt{K} \sqrt{AN} \\

\frac{Y}{AN} = \frac{2 \sqrt{K} \sqrt{AN}}{\sqrt{AN} \sqrt{AN}} \\

Y = 2k^{1/2}"